# Construction and Application of Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model

### v3

@@ Abstract
Knowledge in materials science is widely dispersed across extensive scientific literature, posing significant challenges for efficient discovery and integration of new materials. Traditional methods, often reliant on costly and time-consuming experimental approaches, further complicate rapid innovation. Addressing these challenges, the integration of artificial intelligence with materials science has opened avenues for accelerating the discovery process, though it also demands precise annotation, data extraction, and traceability of information. To tackle these issues, this article introduces the Materials Knowledge Graph (MKG), which utilizes advanced natural language processing techniques, integrated with large language models to extract and systematically organize a decade’s worth of high-quality research into structured triples, contains 162,605 nodes and 731,772 edges. MKG categorizes information into comprehensive labels such as Name, Formula, and Application, structured around a meticulously designed ontology, thus enhancing data usability and integration. By implementing network-based algorithms, MKG not only facilitates efficient link prediction but also significantly reduces reliance on traditional experimental methods. This structured approach not only streamlines materials research but also lays the groundwork for more sophisticated science knowledge graphs.
Introduction
@@
In the contemporary information era, despite notable advancements, the creation and advancement of novel materials still heavily rely on traditional, time-consuming trial-and-error methods intertwined with chemical and physical intuitions [1]. These conventional research approaches significantly impede the life-cycle of high-performance material research. Given the specialization, inherent complexity, and vast knowledge base of material science, researchers focusing on a single direction often struggle to efficiently access and understand material knowledge from multidisciplinary studies. For instance, researchers in solar cell development might not fully comprehend studies related to solid-state batteries or organic light-emitting diodes. Yet, the electronic properties of materials across these different domains are highly related, and researchers in different domains can potentially learn from each other. To accelerate the progress of materials research, there is a pressing need to efficiently integrate knowledge from various disciplines [2]. However, this vital knowledge is scattered across a vast array of over 10 million scientific papers, covering diverse topics and disciplines such as materials preparation and functionalization methods, advanced materials characterization techniques, and the exploration of physical, chemical, and biological properties, along with their applications in fields like electronic devices, clean energy storage and transfer, and mechanical engineering. This fragmentation of knowledge represents a significant barrier to interdisciplinary collaboration and innovation. A critical gap in current research infrastructure is the lack of an effective materials science database that can consolidate this scattered knowledge, facilitating easier access and interdisciplinary integration.
Despite the availability of comprehensive databases such as Scopus, Web of Science, and Crossref that offer structured search capabilities for scientific literature, extracting actionable materials science information from the extensive volumes of published research remains a daunting task. In response, databases like the Materials Project [3], OQMD [4], and NOMAD [5] have been developed to provide a clearer understanding of materials properties. While traditional databases are effective in storing structured data and handling basic queries, they may struggle to capture complex relations and infer new knowledge from the data [6]. Specifically, traditional databases store data in tables with a predefined schema, organizing data into rows and columns, and establish relations between entities using primary and secondary keys. This architecture often requires complex and computationally intensive queries to extract interconnected data, which can hinder the effective mining of knowledge from various scientific findings.
@@
Knowledge graph (KG) is a structured representation of information that models the controlled vocabulary and ontological relations of a topical domain as nodes and edges, enabling complex queries and insights that traditional databases cannot easily provide. The adoption of knowledge graphs offers several advantages, including enhanced data interoperability and visualization [7], the ability to infer new knowledge through relational data analysis or graph neural network [8], and improved data quality and consistency through structured representation [9], [10]. These features make knowledge graphs particularly valuable for integrating diverse information sources, providing a unified view of a domain’s knowledge and more comprehensive contextual relationships, thereby facilitating more informed decision-making and discovery [11]. However, the construction of knowledge graphs in specific fields always requires the participation of a large number of expert [12]. This labor-intensive process not only limits the scalability of KGs but also impacts their performance and timeliness [13]. With the rapid development of natural language processing (NLP), methods for extracting information from unstructured text and constructing knowledge graphs have become more efficient and accurate [14]. For instance, in 2016, the Metallic Materials Knowledge Graph (MMKG) was developed to store materials information from various web data resources [15]. Knowledge graphs tailored to lithium-ion battery cathodes have been constructed, aimed at identifying potential new materials candidates [16]. A user-friendly databases focusing on specific material types, such as Metal-Organic Framework Knowledge Graphs (MOF-KG), have been developed [17]. Recently, a material knowledge graph, MatKG and MatKG2, containing information on material properties, structure, and applications, has been developed [2], [18].
@@
However, these material knowledge graphs face even greater challenges. Firstly, although advancements in NLP technology have reduced the dependency on experts to a certain extent, training data still requires extensive annotation to enhance model accuracy [19]. Secondly, the construction of these knowledge graphs often involves predicting relationships between nodes to form triples, which means the entities represented in the KG are not always based on real instances [20]. This can diminish the authenticity and credibility of the KG. Additionally, this approach makes updating the knowledge graph difficult, as each new node introduced necessitates predicting its relationship with every other node, complicating the maintenance of a dynamic and accurate knowledge graph, especially in advanced fields like material science. Acknowledging these challenges, the emergence of LLMs like GPT and LLaMA represents a breakthrough, offering new solutions to enhance the extraction and credibility of structured information [21], [22]. The fine-tuning technique of LLMs can significantly enhance their performance in specific domain text tasks through training with fewer samples [23],[24]. This means improving the results of NER and RE without requiring a large amount of labor becomes possible and was adopt in our research.
In this paper, we have achieved significant advancements in the development of Materials Knowledge Graph (MKG), a pioneering graph database tailored for the field of materials science. Our contributions are as follows: (1) We propose a method to achieve high accuracy in Named Entity Recognition (NER), Relation Extraction (RE), and Entity Resolution (ER). This method allows for efficient conversion of unstructured text from complex research papers into structured information, effectively capturing crucial information and accelerating the research process in the early stages. This process not only retains the source information for each triplet but also makes updating the KG very convenient. (2) We constructed a dynamic MKG that supports rapid and automated updates, which is invaluable for the aggregation and processing of the growing global data on materials science. This capability ensures that researchers have access to the most current data, enhancing the predictive accuracy as the database expands. (3) We use a well-defined label system that enables easy scaling of our KG and potential integration with other structured databases or KGs like MatKG series and BoschKS [25]. This well-defined labeling not only integrates professionalism into the construction of the knowledge graph but also acts as an effective link between researchers from both academia and industry. (4) We demonstrate a similarity calculation method based on the Jaccard Similarity for materials and applications. This protocol could enable the integration of knowledge and the discovery of novel materials for material applications from the wealth of scientific literature, which can facilitate the design process and inspire innovative ideas for future studies of material science researchers.
Methods
@@
Figure 1(a) presents the elaborate workflow of our study. Initially, through Named Entity Recognition (NER) and Relation Extraction (RE) tasks, we extract structural information pertinent to catalysts, batteries, and solar cells. Subsequent Entity Resolution (ER) and data normalization steps enable the integration of information from these distinct fields into a unified knowledge graph. Specifically, as outlined in Figure 1(b), the process begins with the manual annotation and normalization of the initial training dataset, essential for fine-tuning Large Language Models (LLMs) tailored for NERRE tasks. The inference dataset is then methodically divided into ten batches, facilitating the iterative refinement process that follows. The ER tasks are conducted using advanced NLP technologies, including ChemDataExtractor [26], mat2vec [27], and our expertly curated dictionary. After ER, we selectively enhance the training dataset with high-quality results to improve the model’s performance in subsequent iterations. The knowledge graph is finally constructed using the normalized data from the last iteration. To complete the graph and predict potential material applications, we employ both network-based algorithms and graph embeddings. This methodology provides critical insights and recommendations for researchers in the materials science domain.
Figure 1: (a) Workflow and (b) pipeline of the fine-tuned LLM for knowledge graph tasks.
Data preparation and schema design
@@
Material experts annotated nine distinct categories from the abstracts of 75 research papers, forming the training dataset for the Large Language Models (LLMs). The structure of the data reflects the structure of the KG. As illustrated in Figure 2, the "Material" node serves as the central hub within the graph, linking to nodes that encapsulate its nomenclature, composition, and various attributes. Specifically, the "Material" node comprises sub-nodes such as "Formula," "Name," and "Acronym" for identifying materials. The "Structure/Phase" node describes the material’s physical state, while the "Application" node denotes its practical uses, and the "Property" node details its inherent characteristics. Additionally, the "Descriptor" node provides qualitative information enhancing the contextual understanding of the material.
@@
Furthermore, the "Application" node is extended to include "Property", "Descriptor", and "Domain" nodes, indicating the specific attributes and the broader context of the material’s application. To maintain data provenance and traceability, each node is linked to a "Digital Object Identifier (DOI)" node. This allows for source verification where querying the intersections of "DOI" node connections at both ends of a relation can pinpoint the source article from which the relations was derived, ensuring data integrity within the graph structure.
Figure 2: This schema represents the ontology of MKG (a) All the nodes in MKG connect to all their sources - "DOI". (b) Relations between all labels. (c) Relations between core labels, the more arrows pointing to the label, the lower the priority it has.
@@
Given the inherent complexity of sentences and the variability in terminologies across abstracts, a normalization process was employed after the initial extraction. This normalization ensured uniform representation of entities with similar meanings. For example, terms such as "Lithium-Ion Battery" and "Li-ion batteries" were standardized to "lithium-ion battery", while phrases like "solution casting method", "solvent post-treatment method", and "solution-based deposition" were simplified to "solution-processed", ensuring consistency and facilitating effective training of LLMs. The field of materials encompasses a wide range of areas. At this stage, our priority is to focus on energy materials. We download 150,000 abstracts of peer-reviewed research articles of energy material science includes battery, solar cell and catalyst from the Web of Science. Each abstract was stored in a JSON file format, structured as "DOI - text," facilitating seamless processing and analysis.
LLMs training, evaluation and inference
@@
The training dataset, composed of compiled data, was employed to fine-tune models including LLaMA 7b, LLaMA2 7b [28], and Darwin [29]. Upon obtaining high-quality results from the normalized inference, we iteratively retrain a fine-tuned LLM for infer subsequent batches of data. The models underwent training over 10 epochs with a batch size of 1. Additionally, 60 abstracts were annotated to assess the LLMs" performance. Our evaluation primarily focused on the NER capabilities of the model and preliminarily explored the RE task"s ability to identify potential internal relations among relevant element sets. Moreover, since we standardized entities during the data compilation phase, we also assessed the LLMs" effectiveness in standardizing the entities. Specifically, for NER, RE and ER task, we employ a unified framework for evaluation based on precision, recall, and the F1 score, taking into account the instances of false positives (fp) and false negatives (fn) to quantify the performance.
@@
After evaluation, we choose a fine-tuned LLM, which demonstrates optimal performance in both NER and RE task, to iteratively infer the 150,000 abstracts. The output of inferences are organized in the "DOI - text - response" format. Consequently, the fine-tuned LLM not only extracts entities but also assigns them to appropriate labels, thereby accomplishing NER and RE tasks concurrently. Moreover, every entity and relation identified in the response is traceable, enhancing the integrity and utility of the data.
Entity resolution
The quality of KG is crucial for its credibility, to further check and correct the inference results before graph construction. To ensure the precision of these results, we initially employ ChemDataExtractor to identify chemical formulas and "Name-Acronym" pairs from abstracts. Subsequently, entities recognized by both the Large Language Model (LLM) and ChemDataExtractor are embedded using the mat2vec model. We analyze their similarities to rectify core entities and ensure accurate "Name" to "Acronym" associations. Given the frequent mislabeling of "Name" and "Formula", we progress to the Entity Resolution - Name/Formulas (ER-N/F) phase. Here, we refine the LLM using a specifically curated training set comprising 2,000 accurately labeled entities for binary classification, which is evaluated against an additional set of 200 labels.
For other labels, we implement a Density-Based Spatial Clustering of Applications with Noise [30] algorithm to create the expert dictionary (ER-ED). This algorithm dynamically forms clusters based on vector similarity, without the need to predefine the number of clusters. Each cluster is named by material science experts, leading to the creation of an expert dictionary containing approximately 600 terms related to structures, phases, applications, and more. This dictionary, along with the entities extracted by the LLM, undergoes similarity analysis to standardize and confirm the accuracy of both entities and relations. To elevate the quality of the training dataset, we selectively integrate high-quality data from normalized inference outputs from each iteration into the training set, continuously monitoring and enhancing the performance of the fine-tuned LLM to ensure its efficacy.
Knowledge graph construction

### v5

# neurips_2024.bib

@@ Abstract
Knowledge in materials science is widely dispersed across extensive scientific literature, posing significant challenges to the efficient discovery and integration of new materials. Traditional methods, often reliant on costly and time-consuming experimental approaches, further complicate rapid innovation. Addressing these challenges, the integration of artificial intelligence with materials science has opened avenues for accelerating the discovery process, though it also demands precise annotation, data extraction, and traceability of information. To tackle these issues, this article introduces the Materials Knowledge Graph (MKG), which utilizes advanced natural language processing techniques integrated with large language models to extract and systematically organize a decade’s worth of high-quality research into structured triples, contains 162,605 nodes and 731,772 edges. MKG categorizes information into comprehensive labels such as Name, Formula, and Application, structured around a meticulously designed ontology, thus enhancing data usability and integration. By implementing network-based algorithms, MKG not only facilitates efficient link prediction but also significantly reduces reliance on traditional experimental methods. This structured approach not only streamlines materials research but also lays the groundwork for more sophisticated science knowledge graphs.
11footnotetext: Corresponding authors: shaozhou@greendynamics.com.au; tong@greendynamics.com.au; wenjie.zhang@unsw.edu.au22footnotetext: These authors contributed equally to this work.
1 Introduction
@@
In the contemporary information era, despite notable advancements, the creation and advancement of novel materials still heavily rely on traditional, time-consuming trial-and-error methods intertwined with chemical and physical intuitions. These conventional research approaches significantly impede the life-cycle of high-performance material research. Given the specialization, inherent complexity, and vast knowledge base of material science, researchers focusing on a single direction often struggle to efficiently access and understand material knowledge from multidisciplinary studies. For instance, researchers in solar cell development might not fully comprehend studies related to solid-state batteries or organic light-emitting diodes. Yet, the electronic properties of materials across these different domains are highly related, and researchers in different domains can potentially learn from each other. To accelerate the progress of materials research, there is a pressing need to efficiently integrate knowledge from various disciplines [venugopal2024matkg]. However, this vital knowledge is scattered across a vast array of over 10 million scientific papers, covering diverse topics and disciplines such as materials preparation and functionalization methods, advanced materials characterization techniques, and the exploration of physical, chemical, and biological properties, along with their applications in fields like electronic devices, clean energy storage and transfer, and mechanical engineering. This fragmentation of knowledge represents a significant barrier to interdisciplinary collaboration and innovation. A critical gap in current research infrastructure is the lack of an effective materials science database that can consolidate this scattered knowledge, facilitating easier access and interdisciplinary integration.
Despite the existence of current databases of scientific literature such as Scopus, Web of Science, and Crossref, which offer ways to search for research papers based on specific labels, extracting useful information about material science from the vast ocean of literature remains demanding. To obtain a clearer sense of materials properties, some structured database projects such as Materials Project [jain2013materials], OQMD [saal2013materials], and NOMAD [draxl2019nomad] were developed. However, these databases contain many computational results obtained through techniques like Density Functional Theory (DFT) or Molecular Dynamics (MD) simulations [mrdjenovich2020propnet]. While these computational databases can provide valuable references for predicting and understanding certain materials systems, they often face discrepancies with experimental observations. Therefore, there is an urgent need within the field of materials science for a database grounded in experimental research and practical information.
@@
Knowledge graph (KG) is a structured representation of information that models the controlled vocabulary and ontological relations of a topical domain as nodes and edges, enabling complex queries and insights that traditional databases cannot easily provide. The adoption of knowledge graphs offers several advantages, including enhanced data interoperability, the ability to infer new knowledge through relational data analysis, and improved data quality and consistency through structured representation [Ji], [Zhang2021NeuralSA]. These features make knowledge graphs particularly valuable for integrating diverse information sources and providing a unified view of a domain’s knowledge, thereby facilitating more informed decision-making and discovery [MitchellACM]. However, the construction of knowledge graphs in specific fields always requires the participation of a large number of experts [Zhong]. This labor-intensive process not only limits the scalability of KGs but also impacts their performance and timeliness [PanShirui]. With the rapid development of natural language processing (NLP), methods for extracting information from unstructured text and constructing knowledge graphs have become more efficient and accurate [NLPpromptKG]. For instance, in 2016, the Metallic Materials Knowledge Graph (MMKG) was developed to store materials information from various web data resources [zhang2017mmkg]. Knowledge graphs tailored to lithium-ion battery cathodes have been constructed, aimed at identifying potential new materials candidates [nie2022automating]. User-friendly databases focusing on specific material types, such as Metal-Organic Framework Knowledge Graphs (MOF-KG), have been developed [an2023knowledge]. Recently, a material knowledge graph, MatKG, and MatKG2, containing information on material properties, structure, and applications, has been developed [venugopal2024matkg], [MatKG2].
@@
However, these material knowledge graphs face even greater challenges. Firstly, although advancements in NLP technology have reduced the dependency on experts to a certain extent, training data still requires extensive annotation to enhance model accuracy [Largeannotation]. Secondly, the construction of these knowledge graphs often involves predicting relationships between nodes to form triples, which means the entities represented in the KG are not always based on real instances [sousa2023explainable]. This can diminish the authenticity and credibility of the KG. Additionally, this approach makes updating the knowledge graph difficult, as each new node introduced necessitates predicting its relationship with every other node, complicating the maintenance of a dynamic and accurate knowledge graph, especially in advanced fields like material science. Acknowledging these challenges, the emergence of LLMs like GPT and LLaMA represents a breakthrough, offering new solutions to enhance the zero-shot method [tan2024retrievalenhancedmutationmasteryaugmenting], extraction, and credibility of structured information [GPT], [touvron2023llama]. The fine-tuning technique of LLMs can significantly enhance their performance in specific domain text tasks through training with fewer samples [xie2024creation],[ExtractSI]. This means improving the results of Named Entity Recognition (NER) and Relation Extraction (RE) without requiring a large amount of labor becomes possible and was adopted in our research.
In this paper, we have achieved significant advancements in the development of Materials Knowledge Graph (MKG), a pioneering graph database tailored for the field of materials science. Our contributions are highlighted in three key areas: 1) We propose a method to achieve NER, RE, and entity resolution (ER) with high accuracy. Through this method, we can easily convert unstructured text into triples and retain the source information of each triplet. This method also makes updating KG very convenient. 2) We constructed the first accurate knowledge graph dedicated to materials, where researchers can easily get information about the material by querying the MKG. 3) We use a well-defined label system so our KG can be easily scaled up and potentially combined with other structured databases or KGs. 4) We demonstrate a similarity calculation method based on Jaccard Similarity for materials and applications.
2 Methods
@@
Figure 1 presents the elaborated pipeline of our study. The KG construction part can be divided into there tasks - ontology design, knowledge extraction, and entity resolution (ER). For domain-specific KG, the design of ontology often relies on experts in the field. This work forms an effective ontology by defining and summarizing a small number of papers through LLM. The knowledge extraction task begins with the manual annotation and normalization of the initial training dataset, annotated training set is used to finetune the LLM for NERRE tasks. Simultaneously, the inference dataset is divided into ten batches, facilitating the iterative refinement process that follows. The ER tasks are conducted using advanced NLP technologies, including ChemDataExtractor [CDE], mat2vec [mat2vec], and our expertly curated dictionary. These steps enable the integration of information from these distinct fields into a unified knowledge graph and clean the extracted data. After ER, we selectively enhance the training dataset with high-quality results to improve the model’s performance in subsequent iterations. The knowledge graph is finally constructed using the normalized data from the last iteration. To complete the graph and predict potential material applications, we employ both network-based algorithms and graph embeddings. This methodology provides critical insights and recommendations for researchers in the materials science domain.
Figure 1: Pipeline of the fine-tuned LLM for knowledge graph tasks.
2.1 Data preparation and schema design
@@
Material experts annotated nine distinct categories from the abstracts of 75 research papers, forming the training dataset for the Large Language Models (LLMs). The structure of the data reflects the structure of the KG. This structure summarizes ten articles through LLM and extracts key attributes. As illustrated in Figure 2 (a), the "Formula", "Name" and "Acronym" node serves as the central hub within the graph, linking to nodes that encapsulate its nomenclature, composition, and various attributes. Among these core attributes, the priority of the attribute pointed to by the arrow decreases in sequence. The "Structure/Phase" node describes the material’s physical state, while the "Application" node denotes its practical uses, and the "Property" node details its inherent characteristics. Additionally, the "Descriptor" node provides qualitative information enhancing the contextual understanding of the material.
@@
Furthermore, the "Application" node is extended to include "Property", "Descriptor", and "Domain" nodes, indicating the specific attributes and the broader context of the material’s application. Among them, "Domain" is unique to "Application" and represents the field to which the application belongs. To maintain data provenance and traceability, each node is linked to a "Digital Object Identifier (DOI)" node. This allows for source verification where querying the intersections of "DOI" node connections at both ends of a relation can pinpoint the source article from which the relations were derived, ensuring data integrity within the graph structure. Figure 2 (b) is an example of MKG, from which it can be seen that the core of MKG is to capture the connection and structure between materials and their applications.
Figure 2: This schematic represents the (a) MKG schema and (b) an example of path in MKG between the "Name" node "Copper Indium Gallium Selenide" and "Application" node "Thin Films".
@@
Given the inherent complexity of sentences and the variability in terminologies across abstracts, a normalization process was employed after the initial extraction. This normalization ensured a uniform representation of entities with similar meanings. For example, terms such as "Lithium-Ion Battery" and "Li-ion batteries" were standardized to "lithium-ion battery", while phrases like "solution casting method", "solvent post-treatment method", and "solution-based deposition" were simplified to "solution-processed". All annotated entities underwent this normalization process, ensuring consistency and facilitating effective training of LLMs. The field of materials encompasses a wide range of areas. At this stage, our priority is to focus on energy materials. We downloaded 150,000 abstracts of peer-reviewed research articles on energy material science, including batteries, solar cells, and catalysts, from the Web of Science. Each abstract was stored in a JSON file format, structured as "DOI - text", facilitating seamless processing and analysis.
2.2 LLMs training, evaluation and inference
@@
The training dataset, composed of compiled data, was employed to fine-tune models including LLaMA 7b, LLaMA2 7b [touvron2023llama2], and Darwin [xie2023darwin]. Upon obtaining high-quality results from the normalized inference, we iteratively retrain a fine-tuned LLM to infer subsequent batches of data. The models underwent training over 10 epochs with a batch size of 1. Additionally, 60 abstracts were annotated to assess the LLMs" performance. Our evaluation primarily focused on the NER capabilities of the model and preliminarily explored the RE task’s ability to identify potential internal relations among relevant element sets. Moreover, since we standardized entities during the data compilation phase, we also assessed the LLMs" effectiveness in standardizing the entities. Specifically, for NER, RE, and ER tasks, we employ a unified framework for evaluation based on precision, recall, and the F1 score, taking into account the instances of false positives (FP) and false negatives (fn) to quantify the performance.
@@
After evaluation, we chose a fine-tuned LLM that demonstrates optimal performance in both NER and RE tasks to iteratively infer the 150,000 abstracts. The output of inferences is organized in the "DOI—text—response" format. Consequently, the fine-tuned LLM not only extracts entities but also assigns them appropriate labels, thereby accomplishing NER and RE tasks concurrently. Moreover, every entity and relation identified in the response is traceable, enhancing the integrity and utility of the data.
2.3 Entity resolution
The quality of KG is crucial for its credibility in checking and correcting the inference results before graph construction. To ensure the precision of these results, we initially employed ChemDataExtractor to identify chemical formulas and "Name-Acronym" pairs from abstracts. Subsequently, entities recognized by both the Large Language Model (LLM) and ChemDataExtractor are embedded using the mat2vec model. We analyze their similarities to rectify core entities and ensure accurate "Name" to "Acronym" associations. Through this step, we can make the "Name" and "Formula" different from "Acronym" in MKG. Therefore, we have named this stage "ER-NF/A" ("Entity resolution - Name/Formula and Acronym"). Given the frequent mislabeling of "Name" and "Formula", we progress to the "Entity resolution - Name and Formulas" ("ER-N/F") phase. Here, we refine the LLM using a specifically curated training set comprising 2,000 accurately labeled entities for binary classification, which is evaluated against an additional set of 200 labels.
For other labels, we implement a Density-Based Spatial Clustering of Applications with Noise [ester1996density] algorithm to create the "Entity resolution - expert dictionary" ("ER-ED"). This algorithm dynamically forms clusters based on vector similarity without the need to predefine the number of clusters. Each cluster is named by material science experts, leading to the creation of an expert dictionary containing approximately 600 terms related to structures, phases, applications, and more. This dictionary, along with the entities extracted by the LLM, undergoes similarity analysis to standardize and confirm the accuracy of both entities and relations. To elevate the quality of the training dataset, we selectively integrate high-quality data from normalized inference outputs from each iteration into the training set, continuously monitoring and enhancing the performance of the fine-tuned LLM to ensure its efficacy.
2.4 Knowledge graph construction